OpenStudy (anonymous):

I'm lost on these equations. Could someone explain it to me? -4/2x+3 + 1/x-1 = 1/(2x+3)(x-1)

OpenStudy (anonymous):

I'm wanting to multiply both sides by (2x+3)(x-1)

OpenStudy (anonymous):

This would give me =4x+4 + 2x+3 = 1

OpenStudy (shubhamsrg):

well its -4x and not 4x on LHS

OpenStudy (anonymous):

Yes that was a typo sorry

OpenStudy (apoorvk):

actually, this would give you: -4x+4 + 2x+3 = 1

OpenStudy (apoorvk):

But good other wise, continue!

OpenStudy (anonymous):

Then if continued it would be -2x+7=1

OpenStudy (apoorvk):

yes,

OpenStudy (anonymous):

Then continued after than it would be -2x=-6 so x=3?

OpenStudy (apoorvk):

now, take that 7 to the other side: yeah great workk!!! You 're correct!!!

OpenStudy (anonymous):

Awsome thanks. It just helps me to go step by step with help..

OpenStudy (apoorvk):

No worries ;)